# Generalization Error of the Tilted Empirical Risk

### v2

@@ E.1 Uniform bounds: details for unbounded loss
Now applying $\exp(\gamma\kappa_{u})\leq\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\leq 1$ and $\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\leq\gamma^{2}\kappa_{u}^{2}$ completes the proof. ∎
Corollary 4.4 (Restated).
Under the same assumptions in Theorem 4.3 and a finite hypothesis space, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$, for $\gamma<0$ and with probability at least $(1-\delta)$, the absolute value of the titled generalization error satisfies $\begin{split}\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|&\leq\frac{2\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{(\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta))}{n}}\\ &\qquad-\frac{4\exp(-\gamma\kappa_{u})(\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta))}{3n\gamma(1-\zeta)}-\frac{\gamma}{2}\kappa_{u}^{2}.\end{split}$ (73)
@@
Combining the upper and lower bounds, Theorem 4.2 and Theorem 4.3, we can derive the following bound for a fixed $h\in\mathcal{H}$,
$\begin{split}|\mathrm{gen}_{\gamma}(h,S)|&\leq\frac{2\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{(\log(2/\delta))}{n}}\\ &\qquad-\frac{4\exp(-\gamma\kappa_{u})(\log(2/\delta))}{3n\gamma(1-\zeta)}-\frac{\gamma}{2}\kappa_{u}^{2}.\end{split}$ (74)
Then, using the uniform bound Lemma C.11 completes the proof. ∎
E.2 Information-theoretic Bounds
Assumption E.1 (Bounded second moment).
The learning algorithm $P_{H|S}$, loss function $\ell$, and $\mu$ are such that there is a constant $\kappa_{t}\in\mathbb{R}^{+}$ where the loss function $(H,Z)\mapsto\ell(H,Z)$ satisfies $\mathbb{E}_{P_{H}\otimes\mu}[\ell^{2}(H,Z)]\leq\kappa_{t}^{2}$.
The following results are helpful in deriving the upper and lower bounds under unbounded loss assumptions via the information-theoretic approach.
Proposition E.2.
Given Assumption E.1, the following inequality holds for $\gamma<0$,
$\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\leq\begin{cases}\exp(-\gamma\kappa_{t})\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}},\qquad&\mbox{ if }\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}\\ \frac{-\exp(-\gamma\kappa_{t})}{\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)},\qquad&\mbox{ if }\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}\end{cases}\end{split}$
@@ E.2 Information-theoretic Bounds
For $\gamma<0$ and we use that $0\leq\exp(\gamma\ell(H,\tilde{Z}))\leq 1$ and $\mathrm{Var}(\exp(\gamma\ell(H,\tilde{Z})))\leq\gamma^{2}\mathrm{Var}(\ell(H,\tilde{Z}))\leq\gamma^{2}\kappa_{t}^{2}$. Note that the variable $\exp(\gamma\ell(H,\tilde{Z}))$ is sub-exponential with parameters $(\gamma^{2}\kappa_{t}^{2},1)$ under the distribution $P_{H}\otimes\mu$. Using the approach in (Bu et al., 2020; Aminian et al., 2021a) for the sub-exponential case, we have
$\begin{split}&\Big{|}\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]-\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{|}\\ &\leq\begin{cases}|\gamma|\sqrt{\kappa_{t}^{2}\frac{I(H;S)}{n}}&\quad\mbox{ if}\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}\\ \frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}&\mbox{ if }\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}.\end{cases}\end{split}$ (75)
Therefore, we have for $\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}$,
$\begin{split}\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\leq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}+|\gamma|\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}\Big{)}.\end{split}$ (76)
@@
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{|\gamma|}{\gamma\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}.\end{split}$ (77)
For $\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}$, we have
$\begin{split}&\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\\ &\leq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}+\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)}.\end{split}$ (78)
@@
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{1}{\gamma\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)}.\end{split}$ (79)
As under Assumption E.1 and Lemma C.8, we have $\exp(\gamma\kappa_{t})\leq\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]$, the final result follows. ∎
Corollary E.3.
Given Assumption E.1, the following inequality holds for $\gamma<0$ and some $\zeta\in(0,1)$,
$\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\geq\begin{cases}\frac{-\exp(-\gamma\kappa_{t})}{(1-\zeta)}\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}},\quad&\mbox{ if }\max\big{(}\frac{2I(H;S)}{\gamma^{2}\kappa_{t}^{2}},\frac{\gamma^{2}\kappa_{t}^{2}I(H;S)}{\zeta^{2}\exp(2\gamma\kappa_{t})}\big{)}\leq n\\ \frac{\exp(-\gamma\kappa_{t})}{(1-\zeta)\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)},\quad&\mbox{ if }\min(\frac{4I(H;S)\big{(}4\zeta^{-2}-1\big{)}}{\gamma^{2}\kappa_{t}^{2}},\frac{2I(H;S)}{\gamma^{2}\kappa_{t}^{2}})>n.\end{cases}\end{split}$
@@
Similar to Proposition E.2, we have
$\begin{split}&\Big{|}\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]-\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{|}\\ &\leq\begin{cases}|\gamma|\sqrt{\kappa_{t}^{2}\frac{I(H;S)}{n}}&\quad\mbox{ if}\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}\\ \frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}&\mbox{ if }\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}.\end{cases}\end{split}$ (80)
Therefore, we have for $\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}$,
$\begin{split}\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\geq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}+\gamma\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}\Big{)}.\end{split}$ (81)
@@
$|\gamma|\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}\leq\zeta\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}$ (82)
As $\log(y-x)=\log(y)+\log(1-\frac{x}{y})\geq\log(y)-\frac{x}{y-x}$ for $y>x>0$, then by taking $y=\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}$ and $x=|\gamma|\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}$, so that with (71) we have $y-x\geq(1-\zeta)y>0,$ taking logarithms on both sides of (81) gives that
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{1}{(1-\zeta)\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}},\end{split}$ (83)
where it holds for $n\geq\frac{\gamma^{2}\kappa_{t}^{2}I(H;S)}{\zeta^{2}\exp(2\gamma\kappa_{t})}$. For $\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}$, we have
$\begin{split}&\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\geq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}-\frac{I(H;S)}{n}-\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)}.\end{split}$ (84)

### v3

@@ D.1 Proofs and Details of Uniform bounds under Unbounded Loss
Now applying $\exp(\gamma\kappa_{u})\leq\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\leq 1$ and $\operatorname{Var}(\exp(\gamma\ell(h,\tilde{Z})))\leq 2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{u}^{1+\epsilon}$ completes the proof. ∎
Theorem 3.5 (Restated).
Under the same assumptions in Proposition 3.4 and a finite hypothesis space, then for $n\geq\frac{(4|\gamma|^{1+\epsilon}\kappa_{u}^{1+\epsilon}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$, for $\gamma<0$ and with probability at least $(1-\delta)$, the absolute value of the titled generalization error satisfies $\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|\leq\frac{2\exp(|\gamma|\kappa_{u})}{(1-\zeta)|\gamma|}\sqrt{\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{u}^{1+\epsilon}B(\delta)}{n}}+\frac{4\exp(|\gamma|\kappa_{u})B(\delta)}{3n|\gamma|(1-\zeta)}+|\gamma|^{\epsilon}\kappa_{u}^{1+\epsilon},$ (49) where $B(\delta)=\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)$.
@@
Combining the upper and lower bounds, Proposition 3.3 and Proposition 3.4, we can derive the following bound for a fixed $h\in\mathcal{H}$,
$|\mathrm{gen}_{\gamma}(h,S)|\leq\frac{2\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{u}^{1+\epsilon}(\log(2/\delta))}{n}}+\frac{4\exp(-\gamma\kappa_{u})(\log(2/\delta))}{3n|\gamma|(1-\zeta)}+|\gamma|^{\epsilon}\kappa_{u}^{1+\epsilon}.$ (50)
Then, using the union bound completes the proof. ∎
Lemma 3.7.
The excess risk of the tilted empirical risk satisfies,
@@
It can be proved that,
and
Corollary D.1.
Under the same assumption in Proposition 3.4 and a finite hypothesis space, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ with probability at least $(1-\delta)$ and $\gamma<0$, the excess risk of tilted empirical risk satisfies, $\mathfrak{E}_{\gamma}(\mu)\leq\frac{4\exp(|\gamma|\kappa_{u})}{(1-\zeta)|\gamma|}\sqrt{\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{u}^{1+\epsilon}B(\delta)}{n}}+2|\gamma|^{\epsilon}\kappa_{u}^{1+\epsilon}+\frac{8\exp(|\gamma|\kappa_{u})B(\delta)}{3n|\gamma|(1-\zeta)},$ (51) where $B(\delta)=\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)$ and $\mathfrak{E}_{\gamma}(\mu)$ is defined in (7).
Proof.
Combining Theorem 3.5 with Lemma 3.7 completes the proof. ∎
D.2 Proofs and Details of Information-theoretic Bounds under Unbounded Loss
Proposition D.2.
Under Assumption 3.8, for $\gamma<0$, it holds for some $\epsilon\in(0,1]$ that,
$\begin{split}&0\leq\mathrm{R}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)\leq|\gamma|^{\epsilon}\kappa_{t}^{1+\epsilon},\\ &-|\gamma|^{\epsilon}\kappa_{t}^{1+\epsilon}\leq\mathrm{R}_{\gamma}(H,P_{H,S})-\mathbb{E}_{P_{H,S}}[\widehat{\mathrm{R}}_{\gamma}(H,S)]\leq 0.\end{split}$ (52)
Proof.
The proof follows from Lemma C.9. ∎
Proposition 3.9.
Given Assumption 3.8, the following inequality holds for $\gamma<0$, $\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\leq\begin{cases}\frac{\exp(|\gamma|\kappa_{t})}{|\gamma|}\sqrt{\frac{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{n}},\qquad&\mbox{ if }\frac{I(H;S)}{n}\leq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\\ \frac{\exp(|\gamma|\kappa_{t})}{|\gamma|}\Big{(}\frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\Big{)},\qquad&\mbox{if }\frac{I(H;S)}{n}>\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\end{cases}\end{split}$
Proof.
For $\gamma<0$, we use that $0\leq\exp(\gamma\ell(H,\tilde{Z}))\leq 1$ and $\mathrm{Var}(\exp(\gamma\ell(H,\tilde{Z})))\leq|\gamma|^{1+\epsilon}\mathbb{E}[\ell(H,\tilde{Z})^{1+\epsilon}]\leq|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}$. Note that the variable $\exp(\gamma\ell(H,\tilde{Z}))$ is sub-exponential with parameters $(|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon},1)$ under the distribution $P_{H}\otimes\mu$. Using the approach in (Bu et al., 2020; Aminian et al., 2021a) for the sub-exponential case, we have
$\begin{split}&\Big{|}\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]-\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{|}\\ &\leq\begin{cases}\sqrt{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\frac{I(H;S)}{n}}&\quad\mbox{ if }\frac{I(H;S)}{n}\leq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\\ \frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}&\mbox{ if }\frac{I(H;S)}{n}>\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}.\end{cases}\end{split}$ (53)
Therefore, we have for $\frac{I(H;S)}{n}\leq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}$,
$\begin{split}\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\leq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}+\sqrt{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\frac{I(H;S)}{n}}\Big{)}.\end{split}$ (54)
@@ D.2 Proofs and Details of Information-theoretic Bounds under Unbounded Loss
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{1}{\gamma\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\sqrt{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\frac{I(H;S)}{n}}.\end{split}$ (55)
For $\frac{I(H;S)}{n}>\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}$, we have
$\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\leq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}+\frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\Big{)}.$ (56)
@@
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}-\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])\\ &\geq\frac{1}{\gamma\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\Big{(}\frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\Big{)}.\end{split}$ (57)
As under Assumption 3.8 and Lemma C.8, we have $\exp(\gamma\kappa_{t})\leq\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]$, the final result follows. ∎
Proposition 3.10.
Under Assumption 3.8, there exists $\zeta\in(0,1)$ such that one of the following cases holds, (a) $\zeta\leq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\exp(|\gamma|\kappa_{t})}{2}$ and $n\geq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{\zeta^{2}\exp(2\gamma\kappa_{t})}$, (b) $\zeta>\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\exp(|\gamma|\kappa_{t})}{2}$ and $n\geq\frac{2I(H;S)}{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}$, (c) $\zeta>\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\exp(|\gamma|\kappa_{t})}{2}$ and $\frac{2I(H;S)}{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}>n\geq\frac{2I(H;S)}{2\zeta\exp(\gamma\kappa_{t})-|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}$, Then, the following lower bounds on non-linear expected tilted generalization error hold, $\begin{split}\widehat{\mathrm{gen}}_{\gamma}(H,S)\geq\begin{cases}\frac{-\exp(|\gamma|\kappa_{t})}{|\gamma|(1-\zeta)}\sqrt{\frac{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{n}},\mbox{ \text{if (a) or (b) hold,}}\\ \frac{-\exp(|\gamma|\kappa_{t})}{(1-\zeta)|\gamma|}\Big{(}\frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\Big{)},\mbox{ \text{if (c) holds}}.\end{cases}\end{split}$
@@
Note that, we have,
$\begin{split}&\Big{|}\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]-\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{|}\\ &\leq\begin{cases}\sqrt{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\frac{I(H;S)}{n}}&\quad\mbox{ if }\frac{I(H;S)}{n}\leq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\\ \frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}&\mbox{ if }\frac{I(H;S)}{n}>\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}.\end{cases}\end{split}$ (58)
For $\frac{I(H;S)}{n}\leq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}$, we have,
$\begin{split}\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\geq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}-\sqrt{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\frac{I(H;S)}{n}}\Big{)}.\end{split}$ (59)
@@
$\sqrt{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\frac{I(H;S)}{n}}\leq\zeta\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}$ (60)
As $\log(y-x)=\log(y)+\log(1-\frac{x}{y})\geq\log(y)-\frac{x}{y-x}$ for $y>x>0$, then by taking $y=\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}$ and $x=\sqrt{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\frac{I(H;S)}{n}}$, so that with (47) we have $y-x\geq(1-\zeta)y>0,$ taking logarithms on both sides of (59) gives that
$\begin{split}&\frac{1}{\gamma}\log(\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))])-\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}\\ &\geq\frac{-1}{\gamma(1-\zeta)\mathbb{E}_{P_{H}\otimes\mu}[\exp(\gamma\ell(H,\tilde{Z}))]}\sqrt{\frac{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{n}},\end{split}$ (61)
where it holds for $n\geq\frac{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{\zeta^{2}\exp(2\gamma\kappa_{t})}$. As we also have $\frac{I(H;S)}{n}\leq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}$, we consider the condition $n\geq\max\big{(}\frac{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{\zeta^{2}\exp(2\gamma\kappa_{t})},\frac{2I(H;S)}{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}\big{)}$.
For $\frac{2I(H;S)}{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}>n$, we have
$\begin{split}&\mathbb{E}_{P_{H,S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\geq\Big{(}\mathbb{E}_{P_{H}\otimes\mu}\Big{[}\exp(\gamma\ell(H,\tilde{Z}))\Big{]}-\frac{I(H;S)}{n}-\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\Big{)}.\end{split}$ (62)